# DiffAdapt: Difficulty-Adaptive Reasoning for Token-Efficient LLM Inference

### v1

@@ 1 Introduction
Large Language Models (LLMs) emerged as powerful tools for complex reasoning tasks, spanning mathematical problem-solving (Lewkowycz et al., 2022a; Pan et al., 2024), code generation (Chen et al., 2021), and logical deduction (Wei et al., 2022). A key ingredient in this success is intermediate reasoning steps, often referred to as a “chain of thought” (CoT), before producing a final answer (Yu et al., 2024; Li et al., 2025c; Liu et al., 2024). However, this capability comes at a significant computational cost. Models typically generate a lengthy and elaborate chain of thought for every problem, a process sometimes called test-time scaling (Muennighoff et al., 2025; Chu et al., 2025).
Always generating long traces is fundamentally inefficient. It squanders resources on simple problems, while not necessarily providing sufficient resources for truly complex tasks (Qu et al., 2025; Sui et al., 2025). In this work, we introduce a framework to bridge this gap through systematic empirical analysis and adaptive inference strategy design. We first discover a U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Counterintuitively, models show high uncertainty on simple problems despite achieving high accuracy.
@@
Based on this observation, we introduce DiffAdapt, a three-stage framework that dynamically selects inference strategy rather than applying uniform reasoning budgets to all problems. Our framework operates in three stages: (1) we use a proxy model to generate training data by sampling responses and heuristically labeling them with difficulty-based strategy assignments; (2) we train a lightweight probe on the model’s hidden states to predict problem difficulty; and (3) during inference, the probe dynamically selects the appropriate reasoning strategy (Easy/Normal/Hard) for each question. Compared to the training-free baseline DEER (Yang et al., 2025b), DiffAdapt achieves superior performance with up to 62% token reduction and 18% performance improvement across eight mathematical reasoning benchmarks Lightman et al. (2023); Rein et al. (2024); Wang et al. (2024); He et al. (2024); Cobbe et al. (2021) over five models DeepSeek-AI et al. (2025); Yang et al. (2025a); Liu et al. (2025); Hou et al. (2025).
@@ 2 Related Work
Training-Based Budget Control. Several methods incorporate budget control directly into the model’s training phase. Huang et al. (2025) use reinforcement learning (RL) with difficulty-aware rewards to train a model for adaptive budgeting. Similarly, Cheng et al. (2025) employ a dual-reward system based on Group-Policy Optimization (GRPO) to encourage conciseness. ThinkPrune (Hou et al., 2025) trains long-thinking LLMs via RL with token limits, using iterative pruning rounds to achieve better performance-length tradeoffs. LC-R1 (Cheng et al., 2025) addresses “invalid thinking” through GRPO-based post-training with length and compress rewards, achieving significant sequence length reduction while maintaining performance. TL;DR (Li et al., 2025b) is a training-free method that uses mixed system1 and system2 data to control the reasoning process. AdaCoT (Lou et al., 2025) framed adaptive reasoning as a Pareto optimization problem that seeks to balance model performance with the costs associated with CoT invocation. Thinkless (Fang et al., 2025) is trained under a reinforcement learning paradigm and employs two control tokens, <short> for concise responses and <think> for detailed reasoning.
@@ 6.1 Experimental Setup
Models. We evaluate our framework on five models: three reasoning LLMs (Qwen3-4B (Yang et al., 2025a), DeepSeek-R1-Qwen-7B (DeepSeek-AI et al., 2025), DeepSeek-R1-Llama-8B (DeepSeek-AI et al., 2025)) and three models trained with length control RL training (Nemotron-1.5B (Liu et al., 2025), [^1] https://huggingface.co/nvidia/Nemotron-Research-Reasoning-Qwen-1.5B ThinkPrune-7B (Hou et al., 2025)). The probe uses a simple MLP structure, trained for 100 epochs using the AdamW optimizer with learning rate 1e-3.

### v2

@@ 1 Introduction
Large Language Models (LLMs) emerged as powerful tools for complex reasoning tasks, spanning mathematical problem-solving (Lewkowycz et al., 2022a; Pan et al., 2024), code generation (Chen et al., 2021), and logical deduction (Wei et al., 2022). A key ingredient in this success is intermediate reasoning steps, often referred to as a “chain of thought” (CoT), before producing a final answer (Yu et al., 2024; Li et al., 2025d; Liu et al., 2024). However, this capability comes at a significant computational cost. Models typically generate a lengthy and elaborate chain of thought for every problem, a process sometimes called test-time scaling (Muennighoff et al., 2025; Chu et al., 2025).
Always generating long traces is fundamentally inefficient. It squanders resources on simple problems, while not necessarily providing sufficient resources for truly complex tasks (Qu et al., 2025; Sui et al., 2025; Li et al., 2025b). In this work, we introduce a framework to bridge this gap through systematic empirical analysis and adaptive inference strategy design. We first discover a U-shaped entropy pattern: high entropy on easy problems despite high accuracy, low entropy on problems with medium difficulty, and high entropy on hard problems reflecting uncertainty. Counterintuitively, models show high uncertainty on simple problems despite achieving high accuracy.
@@
Based on this observation, we introduce DiffAdapt, a three-stage framework that dynamically selects inference strategy rather than applying uniform reasoning budgets to all problems. Our framework operates in three stages: (1) we use a proxy model to generate training data by sampling responses and heuristically labeling them with difficulty-based strategy assignments; (2) we train a lightweight probe on the model’s hidden states to predict problem difficulty; and (3) during inference, the probe dynamically selects the appropriate reasoning strategy (Easy/Normal/Hard) for each question. Compared to the training-free baseline DEER (Yang et al., 2025b), DiffAdapt achieves superior performance with up to 62% token reduction and 18% performance improvement across eight mathematical reasoning benchmarks Lightman et al. (2023); Rein et al. (2024); Wang et al. (2024); He et al. (2024); Cobbe et al. (2021) over five models DeepSeek-AI et al. (2025); Yang et al. (2025a); Liu et al. (2025a); Hou et al. (2025).
@@ 2 Related Work
Training-Based Budget Control. Several methods incorporate budget control directly into the model’s training phase. Huang et al. (2025); Shen et al. (2025); Liu et al. (2025b) use reinforcement learning (RL) with difficulty-aware rewards to train a model for adaptive budgeting. Similarly, Cheng et al. (2025) employ a dual-reward system based on Group-Policy Optimization (GRPO) to encourage conciseness. ThinkPrune (Hou et al., 2025) trains long-thinking LLMs via RL with token limits, using iterative pruning rounds to achieve better performance-length tradeoffs. LC-R1 (Cheng et al., 2025) addresses “invalid thinking” through GRPO-based post-training with length and compress rewards, achieving significant sequence length reduction while maintaining performance. TL;DR (Li et al., 2025c) is a training-free method that uses mixed system1 and system2 data to control the reasoning process. AdaCoT (Lou et al., 2025) framed adaptive reasoning as a Pareto optimization problem that seeks to balance model performance with the costs associated with CoT invocation. Thinkless (Fang et al., 2025) is trained under a reinforcement learning paradigm and employs two control tokens, <short> for concise responses and <think> for detailed reasoning.
@@ 6.1 Experimental Setup
Models. We evaluate our framework on five models: three reasoning LLMs (Qwen3-4B (Yang et al., 2025a), DeepSeek-R1-Qwen-7B (DeepSeek-AI et al., 2025), DeepSeek-R1-Llama-8B (DeepSeek-AI et al., 2025)) and three models trained with length control RL training (Nemotron-1.5B (Liu et al., 2025a), [^1] https://huggingface.co/nvidia/Nemotron-Research-Reasoning-Qwen-1.5B ThinkPrune-7B (Hou et al., 2025)). The probe uses a simple MLP structure, trained for 100 epochs using the AdamW optimizer with learning rate 1e-3.